OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of one half. a. y - 4 = one half(x - 1) b. y + 4 = 2(x - 1) c. y + 4 = one half(x - 1) d. y - 4 = 2(x + 1)

OpenStudy (anonymous):

Use the formula \(y-y1=m(x-x1)\) where m is the slope so m=1/2, y1=-4 and x1=1

OpenStudy (anonymous):

So it would be b?

OpenStudy (anonymous):

Close, but in this case m=1/2 so it wouldn't be 2(x-1) it would be \(\frac{1}{2}(x-1)\)

OpenStudy (anonymous):

Oh I see now, thank you:)

OpenStudy (anonymous):

No problem